# Magnetically-Driven Neutron-Rich Ejecta Unleashed: Global 3D Neutrino-GRMHD Simulations of Collapsars Reveal the Conditions for $r$-process Nucleosynthesis

### v1

@@ A.4 Numerical implementation
In our simulations the neutrino-matter interactions are weak (radiative efficiency is $\sim$ few %), thus it is reasonable to separate the evolution of the MHD fluid and neutrino radiation. The evolution algorithm is the following:
@@
    Evolve $T^{\mu}_{\nu}$ and $R^{\mu}_{\nu}$ over the timestep $\Delta t$ (subject to Courant condition) by integrating the fluxes and gravitational source terms explicitly without taking into account neutrino-matter interactions (the source term). For details of computation of fluxes in optically thin and thick regions see A.4.1.
@@
    Using the updated conserved quantities, convert them to primitive quantities. For MHD quantities we used a 2-dimensional Newton-Raphson root-finding method (Noble et al., 2006), and backup entropy-based inversion methods (see Liska et al., 2022, for more details). Neutrino quantities are inverted via the method described in A.4.2.
@@ A.4.1 Fluxes and characteristic speeds
We use Lax-Friedrichs flux for neutrino radiation, for which we need to compute the maximal characteristic wavespeed $a$:
@@
where $\textbf{F}_{L,R},\textbf{U}_{L,R}$ are fluxes and conserved variables computed at the left and right cell interfaces respectively from the reconstructed values of primitive variables using the PPM limiter. We consider both optically thin and thick limits to determine $a$. In the optically thin limit, we take the speed of the radiation characteristic to be $c_{\rm rad}=\pm 1/\sqrt{3}$. In the thick limit, we follow the prescription used in Sądowski et al. (2014); McKinney et al. (2014), where we take
@@ A.4.2 Conserved-to-primitive variables inversion
and primitive quantities are recovered via:
@@ A.4.3 Implicit step
here, $\kappa_{\rm T}=\kappa_{\rm a}+\kappa_{\rm s}$, $w=-n_{\mu}u^{\mu}$, and operating under the assumption that $p^{ij}=P^{ij}/E$ stays constant during this implicit timestep. We end up with a $4\times 4$ matrix inversion per species per timestep. Finally, number density is evolved using the updated values of the primitive quantities.
@@ A.4.4 Neutrino-matter interactions
We use the emission, absorption, and scattering coefficients computed via NuLib (O’Connor, 2015). The table includes $\nu_{\rm e}$ absorption on neutrons, $\bar{\nu}_{\rm e}$ absorption on protons, as well as production of $\nu_{\mu}\bar{\nu}_{\mu}$ and $\nu_{\tau}\bar{\nu}_{\tau}$ pairs from $e^{+}e^{-}$ annihilation and nucleon-nucleon Bremsstrahlung. For electron type neutrinos, we do not include the pair production channels. All inverse reactions are also included, in such a way that the emissivity $\eta$ and absorption opacity $\kappa_{\rm a}$ satisfy Kirchoff’s law $\eta/\kappa_{\rm a}=B_{\nu}$, with $B_{\nu}$ the black-body distribution function of neutrinos in equilibrium with the fluid, integrated over the relevant energy bin. We also include in the table elastic scattering on protons, and neutrons. Emissivity $\eta$, absorption $\kappa_{\rm a}$ and scattering $\kappa_{\rm s}$ opacities from NuLib are functions of density, temperature, electron fraction, and neutrino energy bin. Then, the gray emissivity is $\bar{\eta}=\sum_{b}\eta(E_{b})\Delta E_{b}$ where $E_{b},\Delta E_{b}$ are the energy bin and bin width, summing over all bins. To obtain the gray absorption opacity, we assume Kirchoff’s law $\kappa_{\rm a}(E_{\nu})B_{\nu}(E_{\nu})=\eta(E_{\nu})$, such that the energy averaged opacity is (Foucart et al., 2015):
@@
The scattering opacity is computed using the same expression. In optically thick regions, this prescription is accurate. In the optically thin regions, we use the fact that the cross-sections of the processes used here scale as the square of the average neutrino energy:
@@ B.1 Shadow tests
Figure 6: Lab frame energy density in the radiative shadow tests, where either a single beam (left) or intersecting plane-parallel beams (right) shine on an optically thick sphere (red dashed circle), producing a shadow.
@@
First, we set a single beam of neutrinos at the left boundary, $x=-1,-1\leq y\leq 1$, and solve the problem in 2D (Fig. 6(a)). The grid has a range $x\in(-1,3),y\in(-1.5,1.5)$ with a resolution of $160\times 120$. Initial radiative energy density is $n_{\mu}n_{\nu}R^{\mu}_{\nu}=E=1$ and flux density is $F_{x}=0.999998E$ (corresponding to $\gamma_{R}=2.5\times 10^{5}$). Once we evolve the system, we see a sharp shadow form behind the sphere. This setup is particularly favorable for the M1 scheme due to its simple geometry and single illumination source.
Figure 7: Gaussian pulse propagating in an optically thin medium at different times. The pulse propagates at the speed of light and its peak follows a $\propto 1/r^{2}$ profile (dashed red line).
Next, we set up a test where upper and lower boundaries at $y=-1.5,1.5$ shine planar beams at an inclination of $\theta=\pi/6$ (Fig. 6(b)). The grid has a range $x\in(-6,4),y\in(-1.5,1.5)$ with a resolution of $400\times 120$. Where the beams overlap, we observe the flux directed along the superposition of the directions of two intersecting beams, and energy density around twice-thrice the initial energy densities of the beams, with higher amplification along the midline of intersection.
@@ B.3 Radiative sphere
This test imitates the typical geometry in astrophysical simulations (like a neutron star). We set the sphere of radius $R=1$ where emissivity and absorption opacities are equal $\eta=\kappa_{\rm A}$. This test is performed in 3D and in Cartesian coordinates $x,y,z\in(-2.5,2.5)$ and with resolution of $100^{3}$. Fig. 8 shows the radiation energy density $E=n_{\mu}n_{\nu}R^{\mu}_{\nu}$ in steady state (circles) compared with the analytic solution (solid) for two different values of $\eta=\kappa_{\rm A}$. The analytic solution is obtained from integrating the distribution function (where $\mu=\cos\theta$)
@@ B.4 Single beam propagation in the curved spacetime
As we solve the neutrino transport equations in the vicinity of the BH, we also set up a test to check the propagation of the single beam of neutrinos in curved spacetime. We set the spin parameter to $a_{\rm BH}=0$ and set up a 2D grid in $r\in(1,20)\,{r_{\rm g}},\varphi\in(0,\pi/2)$ at a resolution $150^{2}$. Next we shine beams of neutrinos emanating from the boundary at $\varphi=0$ with $\gamma_{R}=250$, such that $\sqrt{-g}R^{i}_{t}=0$ for $i=r,\theta$ and directed along $\hat{\varphi}$ only. We set the beams at different distances from the BH: $r\in(4.8,5.8)\,{r_{\rm g}}$, $r\in(8.6,9.6)\,{r_{\rm g}}$ and $r\in(16.2,17.1)\,{r_{\rm g}}$. The results are displayed in Fig. 9. We see that beams follow geodesics (red dashed lines) pretty well, and their energy densities undergo some dissipation and widening as they propagate.
B.5 Code benchmarks

### v2

@@ A.4 Numerical implementation
In our simulations the neutrino-matter interactions are weak (radiative efficiency is $\sim$ few percent); thus it is reasonable to separate the evolution of the MHD fluid and neutrino radiation. The evolution algorithm is the following:
@@
    Evolve $T^{\mu}_{\nu}$ and $R^{\mu}_{\nu}$ over the time step $\Delta t$ (subject to Courant condition) by integrating the fluxes and gravitational source terms explicitly without taking into account neutrino-matter interactions (the source term). For details of computation of fluxes in optically thin and thick regions, see A.4.1.
@@
    Using the updated conserved quantities, convert them to primitive quantities. For MHD quantities, we used a two-dimensional Newton-Raphson root-finding method (Noble et al., 2006) and backup entropy-based inversion methods (see Liska et al., 2022, for more details). Neutrino quantities are inverted via the method described in A.4.2.
  4. 4.
    Figure 6 shows the performance of the code on A100 and H100 NVIDIA GPUs. The benchmark is performed using the simulation initialized with the setup of the model cBwon the grid with total size $200\times 200\times 200$ on a single GPU ($2\times 2\times 2$ blocks with $100\times 100\times 100$ cells per block.).
@@ A.4.1 Fluxes and characteristic speeds
We use the Lax-Friedrichs flux for neutrino radiation, for which we need to compute the maximal characteristic wavespeed $a$:
@@
where $\textbf{F}_{L,R},and\textbf{U}_{L,R}$ are fluxes and conserved variables computed at the left and right cell interfaces, respectively, from the reconstructed values of primitive variables using the PPM limiter. We consider both optically thin and thick limits to determine $a$. In the optically thin limit, we take the speed of the radiation characteristic to be $c_{\rm rad}=\pm 1/\sqrt{3}$. In the thick limit, we follow the prescription used in Sądowski et al. (2014) and McKinney et al. (2014), where we take
@@ A.4.2 Conserved-to-primitive variables inversion
and primitive quantities are recovered via
@@ A.4.3 Implicit step
Here $\kappa_{\rm T}=\kappa_{\rm a}+\kappa_{\rm s}$ and $w=-n_{\mu}u^{\mu}$, and operating under the assumption that $p^{ij}=P^{ij}/E$ stays constant during this implicit time step. We end up with a $4\times 4$ matrix inversion per species per time step. Finally, number density is evolved using the updated values of the primitive quantities.
@@ A.4.4 Neutrino-Matter interactions
We use the emission, absorption, and scattering coefficients computed via NuLib (O’Connor, 2015). The table includes $\nu_{\rm e}$ absorption on neutrons, $\bar{\nu}_{\rm e}$ absorption on protons, as well as production of $\nu_{\mu}\bar{\nu}_{\mu}$ and $\nu_{\tau}\bar{\nu}_{\tau}$ pairs from $e^{+}e^{-}$ annihilation and nucleon-nucleon Bremsstrahlung. For electron-type neutrinos, we do not include the pair production channels. All inverse reactions are also included, in such a way that the emissivity $\eta$ and absorption opacity $\kappa_{\rm a}$ satisfy Kirchhoff’s law $\eta/\kappa_{\rm a}=B_{\nu}$, with $B_{\nu}$ the blackbody distribution function of neutrinos in equilibrium with the fluid, integrated over the relevant energy bin. We also include in the table elastic scattering on protons and neutrons. Emissivity $\eta$, absorption $\kappa_{\rm a}$ and scattering $\kappa_{\rm s}$ opacities from NuLib are functions of density, temperature, electron fraction, and neutrino energy bin. Then, the gray emissivity is $\bar{\eta}=\sum_{b}\eta(E_{b})\Delta E_{b}$ where $E_{b}$ and $\Delta E_{b}$ are the energy bin and bin width, summing over all bins. To obtain the gray absorption opacity, we assume Kirchhoff’s law $\kappa_{\rm a}(E_{\nu})B_{\nu}(E_{\nu})=\eta(E_{\nu})$, such that the energy averaged opacity is (Foucart et al., 2015):
@@
The scattering opacity is computed using the same expression. In optically thick regions, this prescription is accurate. In the optically thin regions, we use the fact that the cross sections of the processes used here scale as the square of the average neutrino energy:
@@ B.1 Shadow tests
Figure 7: Lab frame energy density in the radiative shadow tests, where either a single beam (left) or intersecting plane parallel beams (right) shine on an optically thick sphere (red dashed circle), producing a shadow.
@@
First, we set a single beam of neutrinos at the left boundary, $x=-1,-1\leq y\leq 1$, and solve the problem in 2D (Fig. 7). The grid has a range $x\in(-1,3),y\in(-1.5,1.5)$ with a resolution of $160\times 120$. Initial radiative energy density is $n_{\mu}n_{\nu}R^{\mu}_{\nu}=E=1$ and flux density is $F_{x}=0.999998E$ (corresponding to $\gamma_{R}=2.5\times 10^{5}$). Once we evolve the system, we see a sharp shadow form behind the sphere. This setup is particularly favorable for the M1 scheme due to its simple geometry and single illumination source.
Figure 8: Gaussian pulse propagating in an optically thin medium at different times. The pulse propagates at the speed of light, and its peak follows a $\propto 1/r^{2}$ profile (dashed red line).
Next, we set up a test where upper and lower boundaries at $y=-1.5,1.5$ shine planar beams at an inclination of $\theta=\pi/6$ (Fig. 7). The grid has a range $x\in(-6,4),y\in(-1.5,1.5)$ with a resolution of $400\times 120$. Where the beams overlap, we observe the flux directed along the superposition of the directions of two intersecting beams, and energy density around twice to thrice the initial energy densities of the beams, with higher amplification along the midline of intersection.
@@ B.3 Radiative sphere
This test imitates the typical geometry in astrophysical simulations (like an NS). We set the sphere of radius $R=1$ where emissivity and absorption opacities are equal $\eta=\kappa_{\rm A}$. This test is performed in 3D and in Cartesian coordinates $x,y,z\in(-2.5,2.5)$ and with resolution of $100^{3}$. Fig. 9 shows the radiation energy density $E=n_{\mu}n_{\nu}R^{\mu}_{\nu}$ in steady state (circles) compared with the analytic solution (solid) for two different values of $\eta=\kappa_{\rm A}$. The analytic solution is obtained from integrating the distribution function (where $\mu=\cos\theta$)
@@ B.4 Single beam propagation in the curved spacetime
As we solve the neutrino-transport equations in the vicinity of the BH, we also set up a test to check the propagation of the single beam of neutrinos in curved spacetime. We set the spin parameter to $a_{\rm BH}=0$ and set up a 2D grid in $r\in(1,20)\,{r_{\rm g}},\varphi\in(0,\pi/2)$ at a resolution $150^{2}$. Next we shine beams of neutrinos emanating from the boundary at $\varphi=0$ with $\gamma_{R}=250$, such that $\sqrt{-g}R^{i}_{t}=0$ for $i=r,\theta$ and directed along $\hat{\varphi}$ only. We set the beams at different distances from the BH: $r\in(4.8,5.8)\,{r_{\rm g}}$, $r\in(8.6,9.6)\,{r_{\rm g}}$ and $r\in(16.2,17.1)\,{r_{\rm g}}$. The results are displayed in Fig. 10. We see that beams follow geodesics (red dashed lines) pretty well, and their energy densities undergo some dissipation and widening as they propagate.
@@ Appendix C Neutrino field evolution
Figure 11 shows the electron neutrino, antineutrino and heavy lepton neutrino luminosities measured outside the accretion disk, at an extraction radius of $r_{\rm ext}=100\,\,{r_{\rm g}}\approx 600\,\rm{km}$. Figure 12 shows the meridional maps of electron neutrino and antineutrino number densities at different times, similar to Figure 3.
@@
Appendix D Timescales
Figure 13 shows the density-weighted accretion and neutrino-cooling timescales in the disk. Longer cooling timescales reflect decreasing disk densities. Together with the shorter accretion timescales once the MAD state is reached, these effects lead to a decrease in the ignition radius.
Figure 13: Density-weighted average accretion (solid lines) and neutrino cooling (dashed lines) timescales in the disk at a characteristic radius, $r=10\,{r_{\rm g}}$, in all models, as functions of time.